# Euclid Quick Data Release (Q1)

### v2

@@ 1 Introduction
The MER PF realises the merging of photometric information from VIS and NISP, together with information from external photometric surveys: DES (Abbott et al. 2021), CFIS (Ibata et al. 2017), Pan-STARRS (Magnier et al. 2020), WHIGS and WHISHES (see Aihara et al. 2017), Rubin-LSST (Ivezić et al. 2019). The main objective of the MER PF is to provide source catalogues containing object information relevant for Euclid’s science goals.
The primary requirement for which the MER PF has been designed is the production of a image mosaics, and a photometric catalogue of sources. To provide this, it starts from VIS and NISP calibrated frames, processed by OU-VIS (Euclid Collaboration: McCracken et al. 2025) and OU-NIR (Euclid Collaboration: Polenta et al. 2025) respectively, and external surveys’ stacked frames, processed by OU-EXT (Euclid Collaboration: Aussel et al. 2025), together with related ancillary data and PSF models. Separate object detections are performed on both the photometric VIS and NIR (stacking the three NISP bands $J_{\scriptscriptstyle\rm E}$, $H_{\scriptscriptstyle\rm E}$, and $Y_{\scriptscriptstyle\rm E}$) mosaics. A multi-wavelength flux determination is performed on detected objects with suitable photometric techniques. The final product of the MER processing function is a catalogue of sources with unique identification number, multi-wavelength photometric measurements, and ancillary information (morphological estimates, star/galaxy separation, etc.).
The MER PF also provides the required input, as additional information integrated in the final Euclid catalogue, for photometric redshift computation and for spectra extraction for which respectively OU-PHZ (Euclid Collaboration: Tucci et al. 2025) and OU-SIR (Euclid Collaboration: Copin et al. 2025) are in charge as customer OUs of OU-MER. The MER PF catalogues represent also a primary input for shear estimation, which is performed by OU-SHE. The MER products are among the core products delivered in the first Euclid data release, Q1 (Euclid Quick Release Q1 2025; Euclid Collaboration: Aussel et al. 2025), and are used in many science papers.
@@ 2 The Euclid MER software pipeline overview
With the current Q1 release configuration, taking into account that many PEs run in parallel and some of them make use of more than one processing cores, the total execution time is equivalent to approximately 140 core hours. The mosaicing, background-subraction, and PSF calculation PEs account for 13% of the processing time. On average, 10% of the time is spent on the detection and deblending PEs. The large majority (75%) of the processing time is spent running the different photometry methods and on the calculation of the PSF transformation kernels.
@@ 3.2 Mosaicing
For the later measurement steps, specifically the photometry with A-PHOT and T-PHOT (see Sect. 6), we need a multi-band set of co-aligned images, where each pixel represent the same position on the sky. The layout of these co-aligned images is given by setting a pixel scale of $0\overset{\prime\prime}{.}1\,{\rm pixel}^{-1}$ (to properly sample the VIS data) and the dimensions of the Euclid tiles presented in Sect. 3.1.
@@ 3.3 Background subtraction
where $M$ and $\mu$ are the median and mean of the clipped histogram, respectively.
@@ 4.3 Deblending
The deblending process is carried out using a dedicated software component based on ASTErIsM (Tramacere et al. 2016). ASTErIsM is a Python package designed for automatic source detection and classification, leveraging two topometric clustering algorithms: DBSCAN (Ester et al. 1996) for identifying sources and DENCLUE (Hinneburg et al. 1998; Hinneburg & Gabriel 2007) for deblending overlapping sources and identifying galaxy morphological features such as spiral arms. Our internal processing element, with a specific set of configuration parameters for VIS mosaic and NIR stack, inherits the operation of ASTErIsM as a deblender, which is based on the execution of a modified version of the DENCLUE density-based clustering algorithm, where the kernel-density estimation is substituted with a convolution of the image using a given kernel. This algorithm processes, as a function of the configuration parameters, a downsampled version of the segmentation clusters in terms of a distinct spatial scale. The adaptive approach allows for a robust separation of the overlapping sources within each cluster.
@@ 4.4 Merging the VIS and NIR detections
The VIS data have higher resolution ($0\overset{\prime\prime}{.}1$) than NIR ($0\overset{\prime\prime}{.}3$) and a better sampling. This leads to much better results for the detection and deblending. As a consequence, we combine the VIS and NIR datasets such that the merged catalogue and segmentation image contain all VIS objects plus the so-called NIR-only objects, which are the objects that are detect in NIR but not in VIS.
@@ 6 Multi-band photometry
The pipeline provides the direct output of the template-fitting runs, which can be taken as a proxy for the total flux of the objects. Additionally, fluxes measured in the detection band smoothed to the resolution of each other band are also provided; they can be used to obtain an estimate of colours complementary to that obtained with aperture photometry, and used to scale the detection Kron flux to get another total flux value in each band. Typically, one expects $f_{\rm templ,band}\leq f_{\rm tot,band}$, where the latter is the total flux from aperture photometry as described above; this is because the contamination from nearby sources, when present, should be drastically decreased. Figure 5 shows the situation for a random OU-MER tile, and the case of one discrepant galaxy. However, the colors obtained with the two techniques should obviously be reasonably consistent. Figure 6 shows the compared $g$ $H_{\scriptscriptstyle\rm E}$ colors ($g$ is from DECAM) obtained from aperture and template fitting photometry in the same tile, for stars (blue crosses, POINT_LIKE_PROB$>0.7$) and galaxies (red dots, POINT_LIKE_PROB0.7absent0.7$\leq$ 0.7≤ 0.7; see Sect. 7.4), ensuring this is indeed the case. It should be stressed that for Q1 we found that for point-like sources (bright stars in particular), contrary to what was just discussed for extended sources, the total fluxes computed using aperture photometry were found to be more accurate than template-fitting fluxes in NIR bands, most likely due to their higher sensitivity to PSF related issues (see Sect. 9.5). Currently, OU-PHZ uses aperture colors to compute photometric redshifts (see Euclid Collaboration: Tucci et al. 2025).
@@
Figure 6: Difference in $g$-$H_{\scriptscriptstyle\rm E}$ colors obtained from aperture and template fitting photomety for a random OU-MER tile, as a function of $I_{\scriptscriptstyle\rm E}$ magnitude.
@@ 7.4 Star-galaxy separation
As shown in Fig. 9, the S/G separation is only performed in VIS-detected objects. The current rule-of-thumb to select stars in the MER catalogue is: (a) select very bright (even saturated) sources at VIS magnitudes $I_{\scriptscriptstyle\rm E}<17$; and (b) select POINT_LIKE_PROB $>0.96$. Note that the POINT_LIKE_FLAG is defined as POINT_LIKE_PROB $>0.96$ and DET_QUALITY_FLAG $=0$ so this flag can also be used. A more refined object classification, differentiating between star, galaxies and QSO, is performed by PHZ (Euclid Collaboration: Tucci et al. 2025).
@@ 9.1 Astrometry
To assess the accuracy of the MER pipeline astrometry, we measure, for each tile, the mean positional offset with respect to Gaia (Fig. 11). Tiles with larger values of FILLING_FACTOR (Sect. 9.3), with a comparably large offset are located at the border of the field. This is expected since both these offsets originate from systematic residuals in the VIS astrometry (Euclid Collaboration: McCracken et al. 2025), which are not evenly sampled at the border.
Figure 11: Distribution of the astrometric offsets for all tiles in the EDF-S field.
@@ 9.4 PSF
Figure 14: Spatial variation of the MER catalogue-PSF for the VIS band on an example MER tile in the EDF-S field (tile index 102021019). FWHM values are measured fitting a two-dimensional Gaussian model to each propagated PSF stamp. Large-scale variations can be traced back to PSF changes along the VIS focal plane (see Euclid Collaboration: McCracken et al. 2025, for more details). In this example, the sharp FWHM change at RA $\sim$ 65°65°65  ° coincides with the transition between two different Euclid observation.
@@ 9.5 Photometry
Figure 16: Mean magnitude offsets for aperture (A-PHOT) and template fitting (T-PHOT) photometry in EDF-S derived from the results computed by the validation pipeline against cross-matched Gaia sources for the Euclid ($I_{\scriptscriptstyle\rm E}$, $Y_{\scriptscriptstyle\rm E}$, $J_{\scriptscriptstyle\rm E}$ and $H_{\scriptscriptstyle\rm E}$) and EXT-DECAM bands ($g$, $r$, $i$ and $z$). The error bars correspond to the specific standard deviation associated to each mean magnitude offset calculated.
@@ 9.6 Point-like probability
At the tile level we validate POINT_LIKE_PROB against Gaia matches identified in the MER final catalog (See Sect. 17) as shown in Fig. 18 for tile 102021017. There are 1104 objects with Gaia matches and 1070 with a high point-like probability (POINT_LIKE_PROB $>0.96$ ). From these 1070 there are 744 with and 326 without Gaia matches. The not matched objects are concentrated beyond the limiting magnitude of Gaia.

### v3

@@ 1 Introduction
The MER PF realises the merging of photometric information from VIS and NISP, together with information from external photometric surveys: DES (Abbott et al. 2021), CFIS (Ibata et al. 2017), Pan-STARRS (Magnier et al. 2020), WHIGS and WHISHES (see Aihara et al. 2017), and, in the future, Rubin-LSST (Ivezić et al. 2019). The main objective of the MER PF is to provide source catalogues containing object information relevant for Euclid’s science goals.
The primary requirement for which the MER PF has been designed is the production of image mosaics, and a photometric catalogue of sources. To provide this, it starts from VIS and NISP calibrated frames, processed by OU-VIS (Euclid Collaboration: McCracken et al. 2025) and OU-NIR (Euclid Collaboration: Polenta et al. 2025) respectively, and external surveys’ stacked frames, processed by OU-EXT (Euclid Collaboration: Aussel et al. 2025), together with related ancillary data and PSF models. Separate object detections are performed on both the photometric VIS and NIR (stacking the three NISP bands $J_{\scriptscriptstyle\rm E}$, $H_{\scriptscriptstyle\rm E}$, and $Y_{\scriptscriptstyle\rm E}$) mosaics. A multi-wavelength flux determination is performed on detected objects with suitable photometric techniques. The final product of the MER processing function is a catalogue of sources with unique identification number, multi-wavelength photometric measurements, and ancillary information (morphological estimates, star/galaxy separation, etc.).
The MER PF also provides the required input, as additional information integrated in the final Euclid catalogue, for photometric redshift computation and for spectra extraction for which respectively OU-PHZ (Euclid Collaboration: Tucci et al. 2025) and OU-SIR (Euclid Collaboration: Copin et al. 2025) are in charge as customer OUs of OU-MER. The MER PF catalogues represent also a primary input for shear estimation, which is performed by OU-SHE. The MER products are among the core products delivered in the first Euclid data release, Q1 (Euclid Quick Release Q1 2025), and are used in many science papers. The Q1 data release covers $63.1\,\rm deg^{2}$ of the Euclid Deep Fields (EDFs) to the nominal wide-survey depth and contains nearly 30 million objects. For a detailed description of the Q1 data release, we refer the reader to Euclid Collaboration: Aussel et al. (2025).
@@ 2 The Euclid MER software pipeline overview
With the current Q1 release configuration, taking into account that many PEs run in parallel and some of them make use of more than one processing cores, the total execution time is equivalent to approximately 140 core hours per tile. The mosaicing, background-subraction, and PSF calculation PEs account for 13% of the processing time. On average, 10% of the time is spent on the detection and deblending PEs. The large majority (75%) of the processing time is spent running the different photometry methods and on the calculation of the PSF transformation kernels.
@@ 3.2 Mosaicing
For the later measurement steps, specifically the photometry with A-PHOT and T-PHOT (see Sect. 6), we need a multi-band set of co-aligned images, where each pixel represent the same position on the sky. The layout of these co-aligned images is given by setting a pixel scale of $0\overset{\prime\prime}{.}1\,{\rm pixel}^{-1}$ (to facilitate critical Nyquist sampling for the VIS data, which have a resolution of $\sim 0\overset{\prime\prime}{.}2$) and the dimensions of the Euclid tiles presented in Sect. 3.1.
@@ 3.3 Background subtraction
as described in Bertin & Arnouts (1996). $M$ and $\mu$ are the median and mean of the clipped histogram, respectively.
@@ 4.3 Deblending
The deblending process is carried out using a dedicated software component based on ASTErIsM (Tramacere et al. 2016). ASTErIsM is a Python package designed for automatic source detection and classification, leveraging two topometric clustering algorithms: DBSCAN (Ester et al. 1996) for identifying sources and DENCLUE (Hinneburg et al. 1998; Hinneburg & Gabriel 2007) for deblending overlapping sources and identifying galaxy morphological features such as spiral arms. Our internal processing element, with a specific set of configuration parameters for the VIS mosaic and the NIR stack, inherits the operation of ASTErIsM as a deblender, which is based on the execution of a modified version of the DENCLUE density-based clustering algorithm, where the kernel-density estimation is substituted with a convolution of the image using a given kernel. This algorithm processes, as a function of the configuration parameters, a downsampled version of the segmentation clusters in terms of a distinct spatial scale. The adaptive approach allows for a robust separation of the overlapping sources within each cluster.
@@
Figure 5: Two examples of suboptimal deblending results produced by the OU-MER pipeline in problematic scenarios. In the top row, we can see how the segmentation map of a galaxy is highly perturbed by the close proximity of two bright stars. In the bottom row, we can observe how the deblending algorithm struggles when processing galaxies with intricate features such as spiral ones.
To illustrate some of the expected limitations of the pipeline’s deblending step when handling complex cases, we present in Figure 5 two challenging examples, that go beyond the typical merging galaxies and/or dense environment scenario, where the pipeline provides suboptimal results due to the convoluted nature of the situation. In the top row we can observe how the segmentation map of a galaxy located in the vicinity of two bright stars is significantly impacted by a spike originating from one of them. This interference even causes the galaxy’s segmentation map to incorrectly include another nearby source that was not properly deblended. The bottom row showcases an ambiguous case where the arms of a spiral galaxy have been partially overdeblended. This seems to be due to the galaxy’s diffuse nature combined with the presence of internal features or background sources.
@@ 4.4 Merging the VIS and NIR detections
The VIS data have higher resolution ($0\overset{\prime\prime}{.}1$) than NIR ($0\overset{\prime\prime}{.}3$) and a better sampling. This leads to much better results for the detection and deblending. As a consequence, we combine the VIS and NIR datasets such that the merged catalogue and segmentation image contain all VIS objects plus the so-called NIR-only objects, which are the objects that are detected in NIR but not in VIS.
@@ 6 Multi-band photometry
The pipeline provides the direct output of the template-fitting runs, which can be taken as a proxy for the total flux of the objects. Additionally, fluxes measured in the detection band smoothed to the resolution of each other band are also provided; they can be used to obtain an estimate of colours complementary to that obtained with aperture photometry, and used to scale the detection Kron flux to get another total flux value in each band. Typically, one expects $f_{\rm templ,band}\leq f_{\rm tot,band}$, where the latter is the total flux from aperture photometry as described above; this is because the contamination from nearby sources, when present, should be drastically decreased. Figure 6 shows the situation for a random OU-MER tile, and the case of one discrepant galaxy. However, the colors obtained with the two techniques should obviously be reasonably consistent. Figure 7 shows the compared $g$ $H_{\scriptscriptstyle\rm E}$ colors ($g$ is from DECAM) obtained from aperture and template fitting photometry in the same tile, for point-like (blue crosses, POINT_LIKE_PROB$>0.7$) and extended sources (red dots, POINT_LIKE_PROB0.7absent0.7$\leq$ 0.7≤ 0.7; see Sect. 7.4), ensuring this is indeed the case. It should be stressed that for Q1 we found that for point-like sources (bright stars in particular), contrary to what was just discussed for extended sources, the total fluxes computed using aperture photometry were found to be more accurate than template-fitting fluxes in NIR bands, most likely due to their higher sensitivity to PSF related issues (see Sect. 9.5). Currently, OU-PHZ uses aperture colors to compute photometric redshifts (see Euclid Collaboration: Tucci et al. 2025).
@@
[…]
@@ 7.4 Star-galaxy separation
As shown in Fig. 10, the S/G separation is only performed in VIS-detected objects. The current rule-of-thumb to select stars in the MER catalogue is: (a) select very bright (even saturated) sources at VIS magnitudes $I_{\scriptscriptstyle\rm E}<17$; and (b) select POINT_LIKE_PROB $>0.96$. Note that the POINT_LIKE_FLAG is defined as POINT_LIKE_PROB $>0.96$ and bits 1 to 4 of DET_QUALITY_FLAG $=0$ so this flag can also be used.
As a more refined object classification, differentiating between star, galaxies and QSO, is performed by PHZ (Euclid Collaboration: Tucci et al. 2025) using the full photometric coverage and the spectral energy distribution information of each source, we did not use other S/G separation and classification methods (Odewahn et al. 2004; Soumagnac et al. 2015; Slater et al. 2020) at the MER level.
@@ 9.1 Astrometry
As a validation of the positional accuracy of the objects in our final catalog we measure, for each tile, the mean positional offset of the Gaia objects in the catalog with respect to their Gaia reference position. Fig. 12 shows the robust statistics (median plus NMAD) for these offsets as determined for all tiles of the EDF-S field. Tiles with a FILLING_FACTOR (Sect. 9.3) larger than $0.5$ have an offset smaller than $0."01$ in both right ascension and declination, which is the requirement for the absolute accuracy. The ensemble of these tiles in EDF-S has the mean value of $3.4$ and $-0.3\,\mathrm{mas}$ with a rms of $2.2$ and $5.6\,\mathrm{mas}$ in right ascension and declination, respectively. These offsets originate from systematic residuals in the VIS astrometry (Euclid Collaboration: McCracken et al. 2025), which also explains the larger offsets for the tiles at the border of the field with a smaller FILLING_FACTOR, which do not evenly sample these residuals.
Figure 12: Distribution of the astrometric offsets (median and NMAD) for all tiles in the EDF-S field.
@@ 9.4 PSF
Figure 15: Spatial variation of the MER catalogue-PSF for the VIS band on an example MER tile in the EDF-S field (tile index 102021019). FWHM values are measured fitting a two-dimensional Gaussian model to each propagated PSF stamp. White dashed lines indicate the outer spatial footprint of the VIS calibrated frames that intersect the tile. Large-scale variations can be traced back to PSF changes along the VIS focal plane (see Euclid Collaboration: McCracken et al. 2025, for more details).
@@ 9.5 Photometry
Figure 17: Mean magnitude offsets for aperture (A-PHOT) and template fitting (T-PHOT) photometry in EDF-S derived from the results computed by the validation pipeline against cross-matched Gaia sources for the Euclid ($I_{\scriptscriptstyle\rm E}$, $Y_{\scriptscriptstyle\rm E}$, $J_{\scriptscriptstyle\rm E}$ and $H_{\scriptscriptstyle\rm E}$) and EXT-DECAM bands ($g$, $r$, $i$ and $z$). The error bars presented in this plot represent the standard deviations of the specific samples utilized to calculate each mean magnitude offset. This provides valuable insight into the inherent scatter within each one of these individual samples.
@@ 9.6 Point-like probability
At the tile level we validate POINT_LIKE_PROB against Gaia matches identified in the MER final catalog (See Sect. 18) as shown in Fig. 19 for tile 102021017. There are 1104 objects with Gaia matches and 1070 with a high point-like probability (POINT_LIKE_PROB $>0.96$ ). From these 1070 there are 744 with and 326 without Gaia matches. The not matched objects are concentrated beyond the limiting magnitude of Gaia.
@@ Acknowledgements.
ER acknowledge funding by the Agenzia Spaziale Italiana (ASI) under agreement no. 2024-10-HH.0.